# Conditioning LLMs to Generate Code-Switched Text: A Methodology Grounded in Naturally Occurring Data

### v1

@@ 5.2 Error analysis
In order to further explore differences between model performance, we analyse the most common errors made by the CS generation models, both quantitatively and qualitatively. We adapt the methodology presented in Popović (2018), who proposes different typologies for machine translation errors, and extend it to CS generation error analysis. To do so, we randomly select a set of 100 outputs from all models and conduct a detailed examination of the types of errors present in them. This thorough analysis allows us to identify recurring patterns and propose a refined error typology specifically for automatic CS generation. This initial error analysis yields 18 total error categories, which we simplify and group into three main error types: a) CS errors, b) Translation Errors, and c) Format errors. The full error typology, along with detailed descriptions for each error type, is provided in Appendix D, while here we explain the three error categories:
@@
We classify $500$ additional instances ($100$ instances per model, obtained from the same source sentences) into these kind of errors, and show the results in Figure 1. The analysis indicates that the baseline MT system has the worst performance, with nearly twice as many errors as the best models. This is in line with the results of the preference-based evaluation, where MT ranked last. The most frequent errors made by the MT system — word repetition, missing information, and incorrect translations — are particularly problematic, as they often make the output unintelligible or incomplete. Among the pre-trained models, Llama models make the fewest errors overall, with 19 fewer errors on average. Base models of both families struggle mainly with format errors, which make up $50.68\%$ of their errors on average, whereas instruction-tuned models present more meaning-related issues, $53,45\%$. This suggests that the linguistic knowledge of the models degrade when tuned on instructions, a phenomenon that has been observed on other related areas Fu et al. (2024). CS-related mistakes are the least common in all models, accounting for less than $15\%$ of the overall error count. It seems that the models have effectively learned to switch between languages naturally, though they may still be prone to other types of errors.
@@ 6 Automatic Evaluation
Llama3 34.49 81.64 53.17
Llama3 Instruct 33.42 81.77 52.01
@@
Table 6: Results of automatic metrics the EN2CS test set. Best results in bold, second best results underlined.
@@
We perform a quantitative evaluation using traditional metrics used in NLG. To that end, we use BLEU (Papineni et al., 2002), BERTScore (Zhang et al., 2020), and chrF (Popović, 2015), implemented with the evaluate library. All three are reference-based task-agnostic quality metrics that give results between 0-1, based on character-level F-score, n-gram precision and semantic similarity using contextual embeddings [^5] BERTscore has been calculated using the embeddings from the model Bert Base Multilingual Cased. respectively.
The results of the evaluation can be seen in Table 6. The best two models are Llama3, with the highest BLEU and chrF, and Llama3 Instruct, with the highest BERTScore. They are closely followed by less than 2 points in all metrics by Mistral, the third best model overall. Finally, Mistral Instruct is the pre-trained model with the lowest results.
The MT baseline obtains the lowest results overall, which is consistent with the qualitative evaluation described above, where the MT baseline proves to be the worst system by a large margin. However, it is worth noting that not all metrics capture this gap in performance, since, according to BERTScore, there is only a 2 point difference between the MT baseline and Mistral, which is the pre-trained LLM that obtains the worst metrics. It is also interesting to compare these results with the error analysis in Section 5.2. For instance, Mistral Instruct yields low values for BLEU and chrF, in line with its number of translation errors, as depicted in Figure 1. However, automatic metrics fail to capture the fact that CS errors produced by language model based systems are relatively low.
6.1 Correlation With Automatic Metrics
The automatic metrics used in the section before are known to have poor correlation with human judgment in NLG tasks (Sai et al., 2022), and in this section we analyse whether this poor correlation also occurs when evaluating CS generation. To that end, we compare the automatic metrics results with the preference-based scores obtained in Section 5.1.
@@ 6.1 Correlation With Automatic Metrics
We calculate Pearson’s ($\rho$) correlation coefficient at instance-level, using the $550$ instances employed for the error classification and human evaluation (the output of 5 models for $110$ source sentences) [^6] We do not consider the reference CS sentence when calculating the correlations.. Each data point corresponds to the CS output of one particular model for an English source sentence, and we compute the correlation using two values: the score obtained by the model for this instance in the human preference-based evaluation of Section 5.1, and the score it attains if we apply the same strategy using the values of the automatic metrics to determine the winner.
@@
Figure 2: Heatmap of the correlations between human scores and automatic metrics, calculated using the Pearson Correlation Coefficient. The correlations are calculated for all instances, as well as for different subsets of instances, according to the type of errors they exhibit. * indicates statistical significance $(p\leq 0.05)$.
@@
The correlation coefficients are shown in Figure 2. The top part of the figure shows the correlation using all the instances, whereas the bottom part only considers those instances that showed some type of error, according to the error analysis described in Section 5.2. If we consider all the instances, the maximum $\rho$ correlation value is $0.28$, which indicates a low alignment with the scores assigned by automatic metrics and human judgments. The metric with highest correlation is chrF, which seems to confirm previous research that reports a higher correlation with human judgement than other metrics (Popović, 2015).
Regarding the correlations according to the error types, all metrics obtain the highest correlation on errors related to meaning and format, which is expected, as they were originally designed to evaluate MT systems. However, instances with CS errors show the lowest correlation overall, with only chrF showing $\rho$ values above $0.05$. Human evaluators were explicitly asked to never choose instances without CS, but automatic metrics are not sensitive to these nuances, and may assign high scores to instances regardless of whether they contain CS or not.
@@
All in all, these results confirm that several of the most commonly used automatic metrics for NLG have a poor correlation with human judgments when evaluating CS generation. This underscores the need to research more specialized evaluation methods which are designed specifically to capture the nuances of this task.
@@ 7 Conclusion
Our framework consists of back-translating natural code-switched instances (EN-ES) into monolingual English sentences, and using the resulting parallel corpus, dubbed EN2CS, to fine-tune autoregressive models to turn monolingual sentences into CS. This approach has the potential to improve the naturalness of CS generation, as the gold standard CS text is not artificially generated.
We have provided an extensive evaluation of the results of our models. On the one hand, we perform a human evaluation of the outputs of the models with two parts: a human preference based evaluation and an error analysis of a subset of the test sentences, where we find out three types of errors: CS errors, errors in meaning and format errors. On the other, we employ popular NLG metrics to automatically evaluate the results of our models, and calculate the correlation between both types of evaluation.
Both the automatic metrics and the analysis of the outputs of the models, as well as the human evaluation, show that, when fine-tuning the models with LoRA, base models work better than their instruction-tuned counterparts for this task, and that the Llama3 family obtains better scores than the Mistral family. This could be an example where instruction tuning degrades the base model’s linguistic ability Fu et al. (2024) or alternatively could be related to differences in how LoRA fine-tuning affects each model type.
Our analyses show low correlation between human and automatic evaluations, particularly in cases with CS errors. This suggests that current metrics are not adequate for assessing CS generation, which would require more specialized evaluation methods.
We conclude that pre-trained models are able to yield competent results and generate satisfactory outputs, as the error analysis shows the less abundant type of errors are those related to CS, as opposed to problems in fluency, retaining the original meaning of the sentences, and errors in format. Human evaluation shows that their generations are still not on a par with the original instances.
@@
One of the key points of our research is using open-weight LLMs, however, the use of bigger, more powerful restricted-weights LLMs could very plausibly yield better results, even in a zero-shot scenario (Huzaifah et al., 2024).
@@
Appendix A Prompt Tuning
We test the prompts in Table 7, combined with 0-, 1- and 5-shot strategies. The prompts include the instructions explained in different ways, including more or less information.
@@ Appendix A Prompt Tuning
For the few-shot strategies, the prompt includes the following template at the beginning, alongside a set of manually selected examples:
@@ Appendix B Post-edition Guidelines
Appendix C Pairwise Annotation Guidelines
@@
Appendix D Error Typology

### v2

@@ 5.2 Error analysis
In order to further explore differences between model performance, we analyse the most common errors made by the CS generation models, both quantitatively and qualitatively. We extend the machine translation error typology presented in Popović (2018) to CS generation error analysis. To do so, we randomly select a set of 100 outputs from all models and conduct a detailed examination of the types of errors present in them. This thorough analysis allows us to identify recurring patterns and propose a refined error typology specifically for automatic CS generation. This initial error analysis yields 18 total error categories, which we simplify and group into three main error types: a) CS errors, b) Translation Errors, and c) Format errors. The full error typology, along with detailed descriptions for each error type, is provided in Appendix F, while here we explain the three error categories:
@@
We classify $700$ additional instances ($100$ instances per model, obtained from the same source sentences) into these kind of errors, and show the results in Figure 1.
GPT-4ofs makes the fewest errors overall (50), closely followed by NLLB (56). However, 90% of GPT’s errors (45) and 45% on NLLB’s (25) are CS related, indicating that while these systems preserve the meaning of sentences and generate few formatting errors, they often produce entirely monolingual outputs, which is a critical error. In comparison, CS-related mistakes are the least common in fine-tuned LLMs, accounting for less than $15\%$ of the overall error count. This analysis show that fine-tuned LLMs have effectively learned to switch between languages naturally, though they may still be prone to other less critical types of errors.
Among the fine-tuned LLMs, Llama models present 19 fewer errors on average than the Mistral family. Base models of both families struggle mainly with format errors, which make up $50.68\%$ of their errors on average, whereas instruction-tuned models present more meaning-related issues, $53,45\%$. This suggests that the linguistic knowledge of the models degrade when tuned on instructions, a phenomenon that has been observed on other related areas Fu et al. (2024).
@@ 6 Automatic Evaluation
Previous research highlights the challenges of automatically evaluating code-switching (CS) generation, with many existing metrics showing low correlation with human judgments (Srivastava and Singh, 2021; Kuwanto et al., 2024). On the other hand, recent studies show that using LLMs as evaluators or judges can offer a promising alternative to evaluate generation tasks, as they show a higher alignment with human ratings (Chiang and Lee, 2023; Wang et al., 2023). In this section, we present the results of an automatic evaluation with reference-based NLG metrics (BLEU, BERTScore, chrF) and GPT-4o as a judge, as well as their correlation with human preferences.
6.1 Reference-based metrics
@@
Llama3 8B 34.49 81.64 53.17
Llama3 8B Instruct 33.42 81.77 52.01
@@
Table 5: Results of reference-based metrics the EN-CS test set. Best results in bold, second best results underlined.
@@
We report the results of BLEU (Papineni et al., 2002), BERTScore (Zhang et al., 2020), and chrF (Popović, 2015), implemented with the evaluate library. All three are task-agnostic quality metrics that give results between 0-1, based on character-level F-score, n-gram precision and semantic similarity using contextual embeddings [^4] BERTscore has been calculated using the embeddings from the model Bert Base Multilingual Cased. respectively. We compute the metrics for all systems, and include an Identity system that simply returns the provided input as the output.
The results of the evaluation can be seen in Table 5. The best model is NLLB, with the highest scores for the three metrics. It is closely followed by GPT-4ofs, with the second highest BERTScore, and fine-tuned Llama3, with the second highest BLEU and chrF. They are closely followed by Llama3 Instruct and Mistral. The Identity system scores nearly as well as the top-performing models. The strong results from the Identity baseline and few-shot models, which often produce monolinugal outputs, suggest that reference-based metrics assign high scores to models that match only the English part of the reference. This reflects the nature of the task and dataset, and highlights the limitations and artifacts of using reference-based metrics to evaluate code-switched generation.
6.2 GPT as a judge
As a complementary automatic assessment of the outputs of our models, we have implemented a zero-shot pair-wise evaluation using GPT-4o as a judge, mimicking the settings of the human evaluation. Details about the implementation are included in Appendix G. Results are shown in the third and fourth columns of Table 4. GPT shows a strong preference for few-shot models, whereas these models are ranked third- and second-to-last by humans. Based on the error analysis (c.f Section 5.2), few-shot models tend to make many CS-related mistakes, although they are the most fluent. Thus, this disagreement may be caused by the humans adhering to the guidelines and taking the presence of CS as the main criterion, whereas GPT is making decisions based on the style and fluency of the answers. Regarding fine-tuned LLMs, they all share the same relative ranking in both evaluations. Finally, NLLB is ranked last by GPT, while it is the fourth best model overall as ranked by humans. Further research is needed to explain this behaviour, but there may be some stylistic features of the NLLB model’s outputs that are affecting GPT’s preferences.
6.3 Correlation With Human Evaluation
The reference-based metrics used in Section 6.1 are known to have weak correlations with human judgment in NLG tasks (Sai et al., 2022), whereas JudgeLLM-based evaluations seem promising (Chiang and Lee, 2023; Wang et al., 2023). In this section, we compare reference-based metrics and GPT scores with the preference-based scores obtained in Section 5.1.
@@ 6.3 Correlation With Human Evaluation
We calculate Pearson’s ($\rho$) correlation coefficient at instance-level, using the $700$ instances employed for the error classification and human evaluation (the output of 7 models for $100$ source sentences). [^5] We do not consider the reference CS sentences when calculating the correlations. Each data point corresponds to the CS output of one particular model for an English source sentence, and we compute the correlation using two values: the score obtained by the model for this instance in the human preference-based evaluation of Section 5.1, and the score it attains if we apply the same strategy using the values of the reference-based metrics to determine the winner, or, in case of JudgeLLM, the scores given by GPT.
@@
Figure 2: Heatmap of the correlations between human scores and reference-based metrics and scores given by GPT, calculated using the Pearson Correlation Coefficient. The correlations are calculated for all instances, as well as for different subsets of instances, according to the type of errors they exhibit.
@@
The correlation coefficients are shown in Figure 2. The top part of the figure shows the correlation using all the instances, whereas the bottom part only considers those instances that showed some type of error, according to the error analysis described in Section 5.2.
If we consider all the instances, the maximum $\rho$ correlation value with reference-based metrics is $0.22$, which indicates a low alignment with human scores. The metric with highest correlation is chrF, in line with previous research (Popović, 2015). GPT-4ofs shows a $\rho$ of 0.53, which is stronger than reference-based metrics, but still too weak to be regarded as a reliable measure for assessing CS generation.
If we instead consider instances with errors from the human evaluation, there is again a higher correlation between human scores and GPT’s judgements, with a margin of at least 0.28 points. Instances with CS errors show the lowest overall correlation. This likely derives from the fact that human evaluators never prefer an instance without CS as instructed in the guidelines, but reference-based and JudgeLLM-based metrics are not sensitive to these nuances, and may assign high scores to instances regardless of whether they contain CS or not.
@@
All in all, these results confirm that several of the most commonly used reference-based metrics for NLG have a weak correlation with human judgments when evaluating CS generation. This underscores the need to research more specialized evaluation methods designed specifically to capture the nuances of this task that correlate with human judgement.
@@ 7 Conclusion
Our framework consists of back-translating natural code-switched instances (EN-ES) into monolingual English sentences, and using the resulting parallel corpus, dubbed EN-CS, to fine-tune autoregressive models to translate monolingual sentences into CS. This has the advantage of ensuring that the target sentences contain completely natural CS, which has the potential to improve the naturalness of CS generation.
We experiment with fine-tuning base and instruction-tuned LLMs on our dataset using LoRA. For baselines, we include few-shot LLMs (Llama3.3-70B and GPT-4o) and a pretrained NLLB translation system that we also finetune using our dataset. The results indicate that fine-tuned LLMs show higher ranking in a human preference-based evaluation and fewer critical errors than the other baselines, performing better even than proprietary models such as GPT-4o.
We also perform a meta-evaluation of reference-based NLG metrics commonly used for CS evaluation, as well as an LLM judge (GPT-4o). Our analyses show low correlation between human and reference-based evaluations, while the LLM judge achieves moderate correlations. However, particularly in cases with CS errors, no metric is adequate for assessing CS generation. We therefore advocate for more research in specialized evaluation methods.
@@
Appendix A Prompt-tuning for CS-EN translation
[…]
@@ Appendix A Prompt-tuning for CS-EN translation
For the few-shot strategies, the prompt includes the following template at the beginning, alongside a set of manually selected examples that are representative of some phenomena we want to cover in our prompt:
@@ Appendix B Post-edition Guidelines
Appendix C Inter-annotator agreement
Figure 3: Distribution of Adequacy and Fluency scores per annotator.
Figure 3 shows the distribution of the scores given by to annotators on the fluency and adequacy of the instances translated by Command R. Although the inter-annotator agreement shows moderate agreement ($\kappa=0.57$), the distributions between the annotators are very similar to each other.
Appendix D Fine-tuning / Few-shot prompting
Base model
I want to not work and make money. = quiero no trabajar and make money
Instruction-tuned model
user: "I want to not work and make money."
assistant: "quiero no trabajar and make money"
Few-shot prompting
system prompt: "You are a bilingual speaker of English and Spanish. Translate the following English sentence into code-switched text between both languages. Do not add any comments or explanations:"
user: Source example $n$
assistant: Target example $n$
user: "I want to not work and make money."
assistant: …
Table 7: Examples and format of prompts used for finetuning base and instruction-tuned models and for few-shot prompting
In Table 7, we can see the prompt used for a)fine-tuning of base models; b) fine-tuning of instruction-tuned models; and c) 5-shot prompting. For both fine-tuning settings, at inference time the second part of the prompt that contains the target CS sentence is left blank for the model to complete.
Appendix E Pairwise Annotation Guidelines
@@
Appendix F Error Typology
@@
developer: "You are a helpful bilingual system that knows how to code-switch between English and Spanish and how to distinguish natural sentences. Your only job is to judge sentences and output a verdict A, B or T."
user:"Which one of the next two automatically generated sentences with code-switching is more natural? The most important criterion is that the sentences must have code-switching to even be considered eligible. Two sentences can be tied if they are equally wrong.
A: {s1}
B: {s2}
Answer(A/B/T):"
Table 8: Prompt used for GPT to act as a judge.
Appendix G Implementation of GPT as judge
For the implementation of GPT as judge, the developer and user prompts in Table 8 have been used to prompt GPT-4o. To calculate the scores, we first check the answers the directly contain the desired format, "A", "B or "T", which are the most common. For the rest of the outputs that did not follow this format, it is possible to extract the labels using simple regular expressions. Then, the scores are calculated just like the human score: every time a model is voted, it gets $1$ point, and the loser gets 00 points; in case of ties, both models get $0.5$ points each.